OpenStudy (anonymous):

Can someone help me out!!! my birthday is 10/30/2000 Create a quadratic function, f(x), in vertex form. The a should be between 4 and –4, the h will be your birth month, and the k will be your birth day. Write your equation below. Using complete sentences, explain how to convert your birthday function into standard form. Graph your function. Include your graph below. Using complete sentences, explain how to find the average rate of change for f(x) from x = 4 to x = 7.

OpenStudy (freckles):

So what is the vertex form of a quadratic?

OpenStudy (anonymous):

yes what is it

OpenStudy (anonymous):

oh i already got my equation y=4(x-10)^2+30

OpenStudy (freckles):

ok it says to write it in standard form

OpenStudy (freckles):

y=ax^2+bx+c

OpenStudy (anonymous):

well i got y=4x^2-80x+130 but don't know how to explain it

OpenStudy (freckles):

so you need to multiply (x-10)(x-10) first then distribute the 4 in front of it

OpenStudy (freckles):

how did you get 130 there?

OpenStudy (freckles):

(x-10)(x-10)=?

OpenStudy (anonymous):

y=4(x^2-20x+100)+30 y=4x^2-80x+100+30

OpenStudy (freckles):

you didn't distribute the 4 to all the numbers in the ( )

OpenStudy (anonymous):

oh ok

OpenStudy (freckles):

y=4(x^2)-4(20x)+4(100)+30

OpenStudy (freckles):

your first two terms were right y=4x^2-80x+4(100)+30

OpenStudy (freckles):

but your last term should be 4(100)+30

OpenStudy (freckles):

do you see what I mean

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

can u plz order the steps

OpenStudy (freckles):

well we did the first step together which was multiplying (x-10)(x-10) and the distributing the 4 to the product of (x-10) and (x-10)

OpenStudy (freckles):

next step was combining the like terms

OpenStudy (freckles):

to find the average of change from x=a to x=b you must compute the following: \[\frac{f(b)-f(a)}{b-a}\] where \[f(x)=4x^2-80x+430\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

is it f(x)= or y=

OpenStudy (anonymous):

@freckles

OpenStudy (freckles):

y=f(x)

OpenStudy (anonymous):

i mean is it y= 4(x-10)^2+30or f(x)=4(x-10)^2+30

OpenStudy (freckles):

either

OpenStudy (freckles):

y=f(x)

OpenStudy (freckles):

you can call it either it is a name the question said to name it f(x)

OpenStudy (anonymous):

ok thx